OpenStudy (anonymous):

Points V, W, X, Y, and Z are graphed on a number line. The coordinate of point V is –4. What points are at a distance of 9 units from point V? Choose all answers that are correct. A. W = –13 B. X = 5 C. Y = 9 D. Z = 13

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@Doll.face

OpenStudy (anonymous):

@whystudyalone2 @OregonDuck @HELPMEPLZ!! @genny7 @OpenStudier0 @YoungStudier @lloydlloyd732 @lexisimko @ShadowLegendX @pooja195 @Data_LG2 @timo86m

OpenStudy (anonymous):

WILL BE REWARED!

OpenStudy (texaschic101):

ok...your at -4....and you have to go 9 spaces...in both directions -4 + 9 = 5 -4 + (-9) = -4 - 9 = - 13 so your points are 5 and -13......A and B

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Oh my.

OpenStudy (anonymous):

I really could'nt have helped you here, anyway. Math is not exactly my expertise